# nonblind {nonblind}

### v1

@@ 4.1 Experimental Design
We implement a hiring game as part of an online experiment to test subjects’ performance when using recommendation algorithms in a setting where both subjects and algorithms have private information. Specifically, we label 25 hypothetical applicants as good or bad based on the role they are applying for (private information only available to experiment subjects) and their personality type (private information only available to the algorithms), and ask subjects to only hire good applicants. We measure how the accuracy of subjects’ hiring decisions changes in response to different recommendation algorithms (treatments) that vary at the subject level.
@@
For each subject, the experiment proceeds as follows. Subjects are brought to a Qualtrics survey from the recruiting site Prolific and informed they will make hiring decisions on 25 hypothetical applicants, each of which are good or bad types (referred to as the applicant’s ability). They are informed that each applicant’s ability $Y$ is determined by the relationship of the applicant’s personality type $X$ (denoted by letters A–E) to their intended role $U$ (Engineering, Sales, or Communications), as shown in Figure 2. They then answer three comprehension questions about how ability $Y$ is determined and, conditional on answering correctly, are randomized into a treatment. All but one control condition give subjects a recommendation algorithm to assist them with making hiring decisions. Unlike the study subjects, these algorithms can access the applicants’ personality types, but not their role. Subjects are then asked three more comprehension questions about the specific algorithm they will interact with, but continue in the experiment regardless of their answers. Subjects then make 25 hiring decisions with access to the distribution of applicants as represented in Figure 2, the recommendations of their assigned algorithm, and an explanation of that specific recommendation algorithm. An example of a hiring decision is given in Figure 3. When making these hiring decisions, subjects know the applicant’s role, but are unaware of the applicant’s personality type beyond any information conveyed by the recommendation.
@@
Figure 2: Summary of how the information available to the subject (the applicant’s role) and the information of the algorithm (the applicant’s personality type) relate to the applicant’s ability (bad or good). This information is shown to all participants in the experiment and is easily accessible throughout their decision–making process.
@@
Subjects are incentivized to hire good applicants and to not hire bad applicants. Specifically, as compensation for participating in our experiment, subjects receive \$2.04 for their time (estimated to be 12 minutes) and the chance to earn a \$2.00 bonus as a performance incentive. They earn this bonus if their hiring decision on a randomly selected applicant is correct (hire if good, not hire if bad). Note that this induces equal type I and type II losses ($c_{I}=c_{II}$) from the perspective of our subjects, and we analogously evaluate their performance using the percentage of correct decisions taken. Subjects see all 25 applicants in a random order to make their hiring decisions, but don’t learn whether they took correct decisions.
The joint distribution of applicant role, personality type, and ability pictured in Figure 2 is designed to vary subjects’ willingness to Comply with recommendations across applicant roles while allowing the recommendation to provide useful information cases the applicant role is relatively uninformative. When subjects review an applicant for an Engineering role, they know the worker is good. Subjects are much more uncertain for Sales applicants (60% of which are good) and Communications applicants (40% of which are good). The algorithm’s private information (personality type) varies significantly across the cases that subjects are uncertain, thus containing useful information to help them make their decisions. Moreover, both the subject and algorithm’s private information are of equal value in the sense that a rational agent with access to either the role or the personality type would take the correct decision $76\%$ of the time.
@@ 4.2 Treatments
The recommendation algorithm a subject receives represents the treatment in our experiment and is randomized at the subject level. In addition to four treatment arms representing different recommendation algorithms, there is a control arm in which no recommendation is given. These recommendation algorithms vary by the way they map personality types $X$ to recommended actions $R=f^{\textnormal{rec}}(X)\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}},{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}\}$, where the recommendations follow our earlier example and correspond to not hiring, no recommendation, and hiring the applicant, respectively. We now describe the four different algorithms, and link them to the implicit assumptions about compliance and active decisions that motivate their design. Additionally, Table 3 lists the recommendation algorithms that subjects could be assigned to, alongside the recommendations that these algorithms provide across personality types.
@@
[…]
@@
Subjects in this condition did not observe a recommendation algorithm. We use this treatment to collect information on the distribution of unassisted decisions subjects take.
@@
The algorithm in this condition sends a hire recommendation if at least half of the applicants associated with the given personality type are good and sends a not-hire recommendation otherwise. This Predictive algorithm is the optimal decision an algorithm can take by itself in order to minimize overall errors, so it corresponds to $f^{\textnormal{dec}*}$ in the framework of Section 3. It would correspond to the optimal recommendation if compliance was perfect and active decisions were only leading to worse hiring decisions.
@@
These subjects observe a version of the Predictive algorithm that implements a ‘safety check’, causing the recommendation not to send recommendations for some personality types. Specifically, the Triage algorithm only sends the Predictive algorithm’s recommendation of hire if at least half of the Communications applicants associated with the given personality type are good and it only sent the Predictive algorithm’s recommendation of not hire if at least half of the Engineering and Sales applicants associated with the given personality type are bad. As a consequence, this algorithm does not send recommendations for the B and C personality types (which are equivalent). The Triage algorithm is the optimal triage algorithm ($f^{\textnormal{tri}*}$) for a rational agent with the subjects’ private information.
@@
We now summarize the logic behind these algorithms by examining their implicit assumptions about the subjects’ compliance choices and active decisions. Specifically, Table 4 presents each of the algorithms as a solution to a specific combination of assumptions. For compliance choices, we distinguish between perfect compliance (that is, following the recommendation and only making an active decision when no recommendation is given) and selective compliance. By selective compliance, we mean that subjects always take an active decision for non-Engineering candidates, and comply otherwise. For active decisions, we distinguish between completely random decisions and sophisticated decisions. Sophisticated decisions assume that the subject uses their private role information effectively when they take active decisions, that is, hire Engineering and Sales candidates, which is optimal absent recommendations and across each of our treatment arms. For example, if compliance is perfect and decisions are sophisticated, then we assume that subjects follow all recommendations and correctly decide in the absence of a recommendation when they do not receive one.
Using these combinations of behaviors, we can interpret the above treatment arms as optimal recommendation algorithms. First, the Predictive algorithm is optimal with perfect compliance and random active decisions. Second, the assumptions under which the Triage algorithm is optimal are that compliance is perfect and active decisions are sophisticated. In this case, the subject follows the recommendation when it is given, and otherwise relies on their private information to take a decision. Third, the Complementary algorithm is optimal when subjects comply for non-Engineering candidates only, and take random decisions when no recommendation is sent. Fourth, the Complementary Triage recommendations achieve optimal outcomes when subjects do not comply for (and always hire) Engineering candidates, and effectively use their private information to take a sophisticated decision when they do not receive a recommendation. In theory, this treatment arm would achieve the best performance on average, with only one mistake (not hiring a type-E communications candidate).
Compliance $C$
Active decision $D_{\varnothing}$ Perfect Selective
@@
Table 4: Optimal algorithms under different combinations of compliance choices and active decisons.
Although final recommendations only depend on the personality type $X$, we note that constructing these algorithms requires an understanding of the joint distribution of $X$ with the outcomes $Y$ as well as the role $U$. Whether such a joint distribution can be learned in practice depends on the type and quality of data that is available at the time of construction of an algorithm. Here, we focus instead on an analysis of the relative performance of different recommendation algorithms in a setting where we (and the study subjects) know the joint distribution of $(Y,X,U)$.
@@ 4.3 Data and Main Results
Before analyzing the results of the experiment, we describe the data collection process, final sample sizes for each treatment, and basic performance statistics. Among 1675 surveys begun, about 35% failed to answer at least one of the initial comprehension questions and were screened out, leaving us with 1083 subjects that completed our study. Subjects were split approximately equally between all arms except Control, which received half as many observations as the other treatments. Specifically, each of the four treatment arms contained approximately 240 participants who, like the approximately 120 Control subjects who did not get recommendations, each took 25 hiring decisions for a total of around 6,000 decision-responses per treatment and 3,000 for Control. Overall, subjects performed well in the experiment and within the estimated time frame. The average fraction of correct decisions was 78.1% across all subjects. The average duration of the experiment was around twelve minutes. However, both the performance (standard deviation of 13 percentage points) and duration (standard deviation of almost eight minutes) show significant heterogeneity across subjects.
@@
We compare treatments by average subject performance, with the main results reported in Table 5. Subjects in the Predictive treatment, who received recommendations from an algorithm that was trained to predict outcomes, got around 76% of hiring decisions correctly, significantly above the around 70% in the Control group that was not supplied with any recommendations. Subjects in the Triage, Complementary, and Complementary Triage treatments were all designed to provide recommendations that complement human decisions, and on average outperformed subjects in the Predictive treatment. Hence, the data confirms that the algorithm designed to take optimal decisions by itself (given by the Predictive treatment) is not the optimal recommendation algorithm in this case. Across the treatments, subjects assigned to the Compliance algorithm perform best, with over 81% optimal decisions (although this performance is statistically indistinguishable from the Complementary Triage subjects). Overall, the results suggest that subjects tend to take smart compliance choices and frequently (and correctly) override the algorithm for Engineering candidates. At the same time, decisions are inefficiently noisy.
Treatment arm Fraction optimal decisions (%) (SE)
Control 69.5 (0.7)
Predictive 76.2 (0.5)
Complementary 81.6 (0.7)
Triage 78.8 (0.8)
Complementary Triage 80.1 (1.1)
Feasible optimum 96.0
Algorithm by itself 76.0
Random choice 50.0
Table 5: Average subject performance (through fractions of optimal hiring decisions) across treatments. Bootstrap standard errors are given and clustered at the subject level. In addition to subject performance in the experiment, the feasible optimum considers the best theoretically achievable performance by a combination of a recommendation algorithm with three levels and a rational agent, while we also consider the optimal performance of an algorithm by itself as well as the trivial benchmark of a completely random hiring choice based on the flip of a fair coin.
@@
We further compare subject performance to useful reference points in Figure 4 (justified further in Wu et al., 2023). Throughout, we focus on the fraction of cases subjects are taking the optimal decision on an applicant (hire if good, not hire if bad). Within each treatment, we compare subject performance to the performance when following the algorithm perfectly (corresponding to the Triage Effect of Section 3.2), and plot the impact of subject deviations (Response Effect). We also provide the best achievable performance by an algorithm acting alone (to benchmark which treatments were able to outperform algorithmic automation) as well as the performance achievable by a rational agent acting with the subject’s information (to act as a performance upper bound). The decomposition reveals that subject performance tends to improve by deviating from the algorithm (at least in cases other than the Predictive arm, in which case subject performance is equivalent to perfectly following the algorithm). Interestingly, machine-assisted human performance is best for the somewhat simpler Complementary algorithm than for the theoretically optimal Complementary Triage. Overall, we think that the patterns in Figure 4 are in line with good compliance choices, noisy active decisions, and a penalty for overly complex recommendation algorithms.
@@
Our experiment documents complementarity between humans and algorithms when recommendations are designed in the right way. Subjects in the Triage, Complementary, and Complementary Triage treatments all performed significantly better (at a 5% significance level) than humans by themselves (Control) or the algorithm by itself, thus achieving human-AI complementarity, while subjects in the Predictive treatment could only match the performance human or algorithm could (theoretically) achieve by themselves.
@@ 4.4 Additional Analysis and Robustness Checks
To understand performance differences across treatment arms further, we separate subjects’ propensity to hire applicants across roles and personality types in Figure 5, and relate this behavior to the recommendations subjects receive. From the graph, we can identify that subjects’ unassisted decisions mostly match the average quality of candidates (which would be in line with a probability-matching model where subjects match the probability they hire an applicant with the probability they are good). Their compliance choices are generally sophisticated (in that they correctly override the algorithm for Engineering candidates), although somewhat random. Note that our framework predicts that the Complementary algorithm is optimal if subjects make somewhat noisy active decisions but make sophisticated compliance choices. This is in line with our empirical findings, where the Complementary algorithm is the best among those algorithms we tested and likely the optimal recommendation algorithm we could have implemented.

### v2

@@ 4.1 Experimental Setup
We implement a hiring game as part of an online experiment to test subjects’ performance when using recommendation algorithms in a setting where both the subject and algorithm have private information. Specifically, we label 25 hypothetical applicants as good or bad based on the role they are applying for and their personality type, and ask subjects to only hire good applicants. The role is only known to the human subject and the personality type is only known to the algorithm. We measure the change in the accuracy of subjects’ hiring decisions in response to different recommendation algorithms (treatments) that vary at the subject level.
@@
For each subject, the experiment proceeds as follows. Subjects arrive at a Qualtrics survey from the recruiting site Prolific and learn they will make hiring decisions on 25 hypothetical applicants, of which there are good and bad types (referred to as the applicant’s ability). They learn that each applicant’s ability $Y$ is determined by the relationship of the applicant’s personality type $X$ (denoted by letters A–E) to their intended role $U$ (Engineering, Sales, or Communications), as shown in Figure 4. Then they answer three comprehension questions about how ability $Y$ is determined and, conditional on answering correctly, we randomize them into a treatment.
Subjects receive recommendations from algorithms to assist them with making hiring decisions, except for those subjects in the control arm. Unlike the study subjects, these algorithms can access the applicants’ personality types, but not their role. [^3] Subjects then answer three more comprehension questions about the specific algorithm they will interact with, but continue in the experiment regardless of their answers. We leverage these additional questions for robustness checks below. Subjects then make 25 hiring decisions with access to the distribution of applicants as represented in Figure 4, the recommendations of their assigned algorithm, and an explanation of that specific recommendation algorithm. An example of a hiring decision is given in Figure 5. When making these hiring decisions, subjects know the applicant’s role, but are unaware of the applicant’s personality type beyond any information conveyed by the recommendation.
@@
Figure 4: Summary of how the information available to the subject (the applicant’s role) and the information of the algorithm (the applicant’s personality type) relate to the applicant’s ability (bad or good). This information is shown to all participants in the experiment and is easily accessible throughout their decision-making process.
@@
We incentivize subjects to hire good applicants and to not hire bad applicants. Specifically, as compensation for participating in our experiment, subjects receive \$2.04 for their time (estimated to be 12 minutes) and the chance to earn a \$2.00 bonus as a performance incentive. They earn this bonus if their hiring decision on a randomly selected applicant is correct (hire if good, not hire if bad). Note that this induces equal type I and type II losses ($c_{I}=c_{II}$) from the perspective of our subjects, and we analogously evaluate their performance using the percentage of correct decisions taken. Subjects see all 25 applicants in a random order without replacement to make their hiring decisions but don’t learn whether they made correct decisions.
The joint distribution of applicant role, personality type, and ability (pictured in Figure 4) is designed to test subjects’ willingness to Comply with recommendations across applicant roles while allowing the recommendation to provide useful information in cases the applicant role is relatively uninformative. When subjects review an applicant for an Engineering role, they know the worker is good. Subjects are much more uncertain for Sales applicants (60% of which are good) and Communications applicants (40% of which are good). The algorithm’s private information (personality type) varies significantly across the cases in which subjects are uncertain, thus containing useful information to help them make their decisions. Moreover, both the subject and algorithm’s private information are of equal value in that a rational agent with access to either the role or the personality type would take the correct decision $76\%$ of the time.
@@ 4.2 Treatments
The recommendation algorithm a subject receives represents the treatment in our experiment and is randomized at the subject level. In addition to four treatment arms representing different recommendation algorithms, there is a control arm in which no recommendation is given. These recommendation algorithms differ in the way they map personality types $X$ to recommended actions $R=f^{\textnormal{rec}}(X)\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}},{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}\}$, where the recommendations follow our earlier example and correspond to not hiring, no recommendation, and hiring the applicant, respectively. We now describe the four different algorithms and link them to the implicit assumptions about compliance and active decisions that motivate their design. Additionally, Table 3 lists the recommendations that these algorithms provide across personality types.
@@
Table 3: Recommendations that each algorithm sends across applicants’ personality types $X\in\{\textnormal{A},\textnormal{B},\textnormal{C},\textnormal{D},\textnormal{E}\}$. Note that personality types B and C are equivalent for the distribution of outcomes as well as in terms of recommendations given.
@@
Subjects in this condition do not observe a recommendation algorithm. We use this treatment to collect information on the distribution of subjects’ unassisted decisions.
@@
The algorithm in this condition sends a hire recommendation if at least half of the applicants associated with the given personality type are good and sends a not-hire recommendation otherwise. This Predictive algorithm is the optimal decision an algorithm can take by itself to minimize overall errors, so it corresponds to $f^{\textnormal{dec}*}$ in the framework of Section 3. It would correspond to the optimal recommendation if compliance were perfect and active decisions only led to worse hiring decisions.
@@
These subjects observe a version of the Predictive algorithm that implements a ‘safety check’, causing the algorithm not to send recommendations for some personality types. Specifically, the Triage algorithm only sends the Predictive algorithm’s recommendation of hire if at least half of the Communications applicants associated with the given personality type are good and it only sends the Predictive algorithm’s recommendation of not hire if at least half of the Engineering and Sales applicants associated with the given personality type are bad. As a consequence, this algorithm does not send recommendations for the B and C personality types (which are equivalent). The Triage algorithm is the optimal triage algorithm ($f^{\textnormal{tri}*}$) for a rational agent with the subjects’ private information.
@@
We summarize the logic behind these algorithms by examining their implicit assumptions about the subjects’ compliance choices and active decisions. Specifically, Table 4 presents each algorithm as a solution to a specific combination of assumptions. For compliance choices, we distinguish between perfect compliance (that is, following the recommendation and only making an active decision when no recommendation is given) and selective compliance. By selective compliance, we mean that subjects always make an active decision for Engineering candidates and comply otherwise. For active decisions, we distinguish between completely random decisions and sophisticated decisions. Sophisticated decisions assume that the subject uses their private role information effectively when they take active decisions, that is, only hiring Engineering and Sales candidates, which is optimal absent recommendations and across each of our treatment arms. For example, if compliance is perfect and decisions are sophisticated, then we assume that subjects follow all recommendations and correctly decide in the absence of a recommendation when they do not receive one.
Using these combinations of behaviors, we can interpret the above treatment arms as optimal recommendation algorithms. First, the Predictive algorithm is optimal with perfect compliance and random active decisions. Second, the assumptions under which the Triage algorithm is optimal are that compliance is perfect and active decisions are sophisticated. In this case, the subject follows the recommendation when it is given and otherwise relies on their private information to decide. Third, the Complementary algorithm is optimal when subjects comply for non-Engineering candidates only, and make random decisions when no recommendation is sent. Fourth, the Complementary Triage recommendations achieve optimal outcomes when subjects do not comply for (and always hire) Engineering candidates, and effectively use their private information to make a sophisticated decision when they do not receive a recommendation. If the human were replaced by a rational agent, this treatment arm would achieve the best performance on average, with only one mistake (not hiring a type-E communications candidate).
Compliance $C_{U}$
Active decision $D^{\varnothing}_{U}$ Perfect Selective
@@
Table 4: Optimal algorithms under different combinations of compliance choices and active decisions.
Although final recommendations only depend on the personality type $X$, we note that constructing these algorithms requires an understanding of the joint distribution of $X$ with the outcomes $Y$ as well as the role $U$. Whether such a joint distribution can be learned in practice depends on the type and quality of data that is available at the time of the construction of the algorithm. Here, we focus instead on an analysis of the relative performance of different recommendation algorithms in a setting where we (and the study subjects) know the joint distribution of $(Y,X,U)$. In Section 5, we will return to the question of how to design recommendations when the distribution of $U$ remains unknown and our only insight into the humans’ behavior is through their unassisted decisions ($D^{0}_{U}$).
@@ 4.3 Data and Main Results
Before analyzing the results of the experiment, we describe the data collection process, final sample sizes for each treatment, and basic performance statistics. Among 1675 surveys begun, about 35% failed to answer at least one of the initial comprehension questions and were screened out, leaving us with 1083 subjects that completed our study. Subjects were split approximately equally between all arms except Control, which received half as many observations as the other treatments. Specifically, each of the four treatment arms contained approximately 240 participants who, like the approximately 120 Control subjects who did not get recommendations, each took 25 hiring decisions for a total of around 6,000 decision responses per treatment and 3,000 for Control. Overall, the subjects performed well in the experiment and within the estimated time frame. The subjects of our experiment made the correct decision 78.1% of the time. The average duration of the experiment was around twelve minutes. However, both the performance (standard deviation of 13 percentage points) and duration (standard deviation of almost eight minutes) show significant heterogeneity across subjects.
@@
We compare treatments by average subject performance, with the main results reported in Table 5. Subjects in the Predictive treatment, who received recommendations from an algorithm that was trained to predict outcomes, got around 76% of hiring decisions correct, significantly above the around 70% in the Control group that was not supplied with any recommendations. The Triage, Complementary, and Complementary Triage treatments were all designed to provide recommendations that complement human decisions, and subjects in these groups on average outperformed subjects in the Predictive treatment. Hence, the data confirms that the algorithm designed to take optimal decisions by itself (given by the Predictive treatment) is not the optimal recommendation algorithm in this case. Across the treatments, subjects assigned to the Complementary algorithm perform best, with over 81% optimal decisions (although this performance is statistically indistinguishable from the Complementary Triage subjects). Overall, the results suggest that subjects tend to make smart compliance choices and frequently (and correctly) override the algorithm for Engineering candidates. At the same time, decisions are inefficiently noisy.
Treatment arm Optimal (%) (SE) Hire (%) (SE) Deviated (%) (SE)
Control 69.5 (0.7) 67.0 (1.1) N/A
Predictive 76.2 (0.5) 75.3 (0.7) 19.7 (0.9)
Complementary 81.6 (0.7) 66.9 (0.6) 25.6 (0.8)
Triage 78.8 (0.8) 65.9 (0.6) 23.5 (1.0)
Complementary Triage 80.1 (1.1) 60.9 (0.8) 28.2 (1.0)
Table 5: For each treatment we list the percentage of subjects decisions that were (i) optimal, (ii) hire, or (iii) deviating from the recommendation. Bootstrap standard errors are clustered at the subject level.
@@
We compare subject performance to useful reference points in Figure 6 (justified further in Wu et al., 2023). Throughout, we focus on the fraction of cases in which subjects make the optimal decision on an applicant (hire if good, not hire if bad). Within each treatment, we compare subject performance to the performance when following the algorithm perfectly (corresponding to the Triage Effect of Section 3.2), and plot the impact of subject deviations (Response Effect). We also provide the best achievable performance by an algorithm acting alone (to benchmark which treatments were able to outperform algorithmic automation) as well as the performance achievable by a rational agent acting with the subject’s information (to act as a performance upper bound). The decomposition reveals that subject performance tends to improve by deviating from the algorithm (at least in cases other than the Predictive arm, in which case subject performance is equivalent to perfectly following the algorithm). Interestingly, machine-assisted human performance is better for the somewhat simpler Complementary algorithm than for the theoretically optimal Complementary Triage. Overall, we think that the patterns in Figure 6 are in line with good compliance choices, noisy active decisions, and a penalty for overly complex recommendation algorithms.
@@
Our experiment documents complementarity between humans and algorithms when recommendations are designed well. Subjects in the Triage, Complementary, and Complementary Triage treatments all performed significantly better (at the 5% significance level) than humans by themselves (Control) or the algorithm by itself, thus achieving human–AI complementarity, while subjects in the Predictive treatment could only match the performance humans or algorithms could (theoretically) achieve by themselves.
@@ 4.4 Additional Analysis and Robustness Checks
To understand performance differences across treatment arms further, we separate subjects’ propensity to hire applicants across roles and personality types in Figure 7, and relate this behavior to the recommendations subjects receive. From the figure, we can identify that subjects’ unassisted decisions mostly match the average quality of candidates (which would be in line with a probability-matching model where subjects match the probability they hire an applicant with the probability they are good). Their compliance choices are generally sophisticated (in that they correctly override the algorithm for Engineering candidates), although somewhat random. Note that our framework predicts that the Complementary algorithm is optimal if subjects make somewhat noisy active decisions but make selective compliance choices. This is in line with our empirical findings, where the Complementary algorithm is the best among those algorithms we tested and likely the optimal recommendation algorithm we could have implemented.